# Funny-Valen-Tine: Planning Solution Distribution Enhances Machine Abstract Reasoning Ability

### v2

@@ IV-B Valen targeting Bongard-Logo problem
We consider Valen as a good baseline model for dealing with image progressive pattern reasoning problems, but it can also handle image clustering reasoning problems such as Bongard-Logo. To adapt Valen to the Bongard-Logo problem, we need to transform this problem accordingly.
In a particular instance of the Bongard-Logo problem, we use $x_{i}$ to represent all the images. Specifically, the images in the primary group of this instance are denoted by $\{x_{i}|i\in[1,6]\}$, while those in the auxiliary group are represented by $\{x_{i}|i\in[8,13]\}$. Furthermore, $x_{7}$ represents the test sample that should be assigned to the primary group, and $x_{14}$ represents the test sample that should be assigned to the auxiliary group.
This paper proposes that we can adopt different perspectives to view image clustering reasoning problems, such as the Bongard-Logo problem. Specifically, we can treat them as an option evaluation problem akin to the RPM problem. To elaborate, the images in the primary group $\{x_{i}|i\in[1,6]\}$ can be regarded as a problem statement similar to that in the RPM problem, while the test images and images in the auxiliary group $\{x_{i}|i\in[7,14]\}$ can be seen as an option pool anaLogous to the one in the RPM problem. This transformation process can be visualized through Figure 8.
@@
After such transformation, Valen can be effectively applied to image clustering reasoning problems like the Bongard-Logo problem.
It is noteworthy that when migrating Valen to Bongard-Logo, some adjustments are still necessary to accommodate the data characteristics of Bongard-Logo. Specifically, due to the scarcity of training instances in the Bongard-Logo problem, the ViT-based image representation extractor may struggle to perform effectively[17, 28]. Furthermore, although we transform the Bongard-Logo problem into an RPM-like format, it is unnecessary to analyze the progressive patterns among Bongard-Logo images, as such patterns do not exist in clustering reasoning problems like Bongard-Logo.
Therefore, we have redesigned the image representation extractor, originally illustrated in Figure 4, into a convolutional network concatenated with a Transformer-Encoder. The new extractor design can be visualized in Figure 9.
@@
In the figure 9, we can observe that when a Bongard-Logo image is processed by the new Valen’s feature extractor, it is first encoded into a feature map by ResNet18. Subsequently, this feature map is decomposed into multiple tokens and fed into a standard Transformer-Encoder to obtain multi-viewpoint representations. Such a design allows for the combination of CNN’s ability to analyze a smaller number of samples[1, 2, 3, 17] and ViT’s ability to analyze samples from multiple viewpoint[28, 24]. This design has been applied in previous work[24].
@@
In the figure10, we can see that we have abandoned the use of MLP to extract the row and column progressive information from the incomplete matrix in Bongard-Logo, because the transformed Bongard-Logo problem only contains clustering patterns and not progressive patterns. Instead, we directly feed the representations of all images within the incomplete matrix of Bongard-Logo as tokens into the Transformer-Encoder to extract $M$ clustering patterns of Bongard-Logo. When Valen processes Bongard-Logo problems, $M$ is set to 6.
@@ V Tine, a Method for Advancing the Solution Distribution Boundary
Current visual abstract reasoning problem solvers, including Valen, mostly follow a pattern: solvers are keen on assigning probabilities to each option of reasoning problems. The loss functions designed for these solvers are diverse, and the backbone network designs of these solvers vary greatly, but most of them adhere to this paradigm. We name this type of solver as “probability-highlighting solver”.
@@
Therefore, we can interpret these solvers as a conditional probability distributions obeyed by abstract reasoning problems, or understand them as providing distributions of solutions to abstract reasoning problems. If we use RPM solvers as examples, then these solvers can be abstracted as a conditional probability distribution, denoted as $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$. (Accordingly, when Valen is applied to Bongard-Logo problems, it can be abstracted as a conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{6};\theta)$). In this context, the two conditions of the conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ are the statement of the reasoning problem $\{x_{i}\}_{i=1}^{8}$ and the optimizable parameters $\theta$ of the solver. Here, the $x_{\alpha}$ refers to the option that is pending verification. This conditional distribution implies that, when given a problem statement $\{x_{i}\}_{i=1}^{8}$ and an option $x_{\alpha}$, it is possible to calculate the probability of $x_{\alpha}$ being a solution to $\{x_{i}\}_{i=1}^{8}$.
@@
The optimization of probability-highlighting solver can be regarded as optimizing the conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$. A wide range of loss functions can be used for optimization, but they all share the same objective: to assign high probabilities to the primary samples in the probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ and low probabilities to the auxiliary options. Such an optimization objective makes us realize that the optimization process of the probability-highlighting solver is not aimed at fitting the distribution of correct solutions to visual abstract reasoning problems, but rather is dedicated to fitting a distribution jointly delimited by primary and auxiliary samples.
@@
To simplify the explanation of our findings, we assume that the conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ follows a Gaussian distribution. Consequently, we can explicitly identify the distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ as the distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\mu_{\theta},\sigma_{\theta})$. Then, the negative logarithm of the probability density function of $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\mu_{\theta},\sigma_{\theta})$ can be expressed as follows:
$-\log(P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\mu_{\theta},\sigma_{\theta}))=\frac{{(x_{a}-\mu_{\theta})^{2}}}{{2\sigma^{2}_{\theta}}}-\log(|\,\sigma_{\theta}|\,)$ (1)
Evidently, we have omitted the constant term $-1/2\log(2\pi)$. In this formula, $\mu_{\theta}$ and $\sigma_{\theta}$ respectively represent the mean and standard deviation of $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$. Since we have defined the form of distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ as a Gaussian distribution, we consequently abstract the optimizable parameters $\theta$ into optimizable mean $\mu_{\theta}$ and standard deviation $\sigma_{\theta}$.
@@
Optimizing the distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\mu_{\theta},\sigma_{\theta})$ using the loss function is equivalent to optimizing the values of $\mu_{\theta}$ and $\sigma_{\theta}$. Our optimization objective is that when primary samples (the correct options in an RPM problem, or the samples within the main group in a Bongard-Logo problem) are plugged into formula (1), we expect to obtain a smaller value. Conversely, when auxiliary samples (the incorrect options in an RPM problem, or the samples within the auxiliary group in a Bongard-Logo problem) are plugged into formula (1), we hope that the formula returns a larger value.
@@
Evidently, constraining the return value of the formula to achieve a minimum is equivalent to optimizing the value of $\mu_{\theta}$ in the formula. Conversely, constraining the return value to achieve a maximum is equivalent to minimizing the value of $\sigma_{\theta}$ in the formula. In other words, the probability-highlighting solver positions the mean of the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ for the reasoning problem through primary samples, while the standard deviation of the solution distribution is constrained by auxiliary samples.
@@
It appears to be a discouraging reality that the learning objective of a probability-highlighting solver is not to determine the distribution of the correct solutions to reasoning problems, but rather to approximate a distribution that is delimited by both primary and auxiliary samples within those problems. This suggests that the reasoning accuracy of the probability-highlighting solver is constrained. The standard deviation of the distribution delimited by the auxiliary samples is necessarily greater than or equal to the standard deviation of the correct solution distribution. In other words, the distribution delimited by both primary and auxiliary samples is often much coarser than the distribution of correct solutions. Aiming to approximate the former as the optimization objective is akin to “aiming for mediocrity and achieving even less”. This discouraging reality can be visualized in Figure 11. Note that the figure 11 is intended to schematically represent our analysis results, rather than to depict the actual form of the solution distribution or the correct solution distribution.
@@
This paper argues that to further improve the reasoning accuracy of Valen in solving RPM problems and Bongard-Logo problems, it is necessary to provide Valen with more judicious auxiliary samples. By “more judicious auxiliary samples,” we refer to those that can better delimit the solution distribution provided by Valen. These more judicious auxiliary samples have the potential to narrow down the standard deviation of the solution distribution. The visualization of these more judicious auxiliary samples can be found in Figure 12.
@@
As seen in the figure 12, the judicious auxiliary samples we desire are those that appear more authentic or closer to the primary samples from Valen’s perspective, compared to the original auxiliary samples provided in the reasoning problems. Since the property of being “judicious” is evaluated from Valen’s perspective, whether an auxiliary sample is judicious or not, and the extent of its judiciousness, lacks interpretability from a human perspective. This implies that, relative to the primary samples, samples with smaller attribute shifts or fewer shifted attributes may not necessarily be considered more judicious from Valen’s perspective.
Therefore, this paper argues that we can adopt an adversarial learning strategy to provide Valen with auxiliary samples that are more judicious from Valen’s perspective. Given this insight, we design Tine (Technique for edge Identification of solution distribution based on adversarial Networks) method. Specifically, we intend to emulate adversarial generation techniques similar to GAN[7] or WGAN[33] to generate more judicious samples for Valen. However, unlike GAN or WGAN, whose objective is to train robust and high-performance generators, our utilization of adversarial learning aims to refine Valen, which serves as the discriminator in this context.
Our objective is to design an auxiliary sample generator for Valen. The task of this generator is clear: to produce auxiliary samples that can achieve higher scores (probability) in Valen compared to the original auxiliary samples provided in the reasoning tasks. Conversely, Valen’s role is to classify these auxiliary samples as low-scoring ones. During the training process, the parameters of Valen and this generator are updated alternately and exclusively, with their performance gradually refined through adversarial learning, which is a common process in adversarial learning.
Given that our pursuit is to generate more judicious auxiliary samples from Valen’s perspective, we are not concerned with their interpretability, attribute geometry, or whether they strictly conform to the nature of reasoning problems. Therefore, we believe that instead of forcing the generator to produce image-based auxiliary samples, which would incur greater computational overhead and optimization difficulties, it is more effective to generate representations of judicious auxiliary samples. By solely generating multi-viewpoint representations of reasoning problems, the generator’s structure becomes simpler and easier to optimize. Consequently, our generator is designed as shown in Figure 13. The figure 13 depicts Valen, which is tailored for addressing the RPM problem, and its accompanying auxiliary sample representation generator. Since we have already transformed the Bongard-Logo problem into an RPM-like problem, we will not further elaborate on the structure of the generator for the Bongard-Logo version.
@@
As depicted in Figure 13, the left gray dashed box encloses the previously designed Valen, while the right gray dashed box encloses the generator designed for Valen. The figure 13 shows that the backbone of this generator is a standard Transformer-Encoder, which generates representations of more judicious auxiliary samples based on the representations of the abstract reasoning problem statement $\{r_{i}|i\in[1,8]\}_{v}$, the progressive patterns $\{P_{9m}|m\in[1,M]\}_{v}$ of the statement, and $L$ learnable vectors. Both the representations of the statement and their progressive patterns are provided by Valen, while the $L$ learnable vectors require additional establishment for the generator. It is noteworthy that when dealing with PGM problems in RPM tasks, we have designed special positional encoding for the representations of the problem statement. This positional encoding can be specifically attached to incomplete matrices with missing option positions and is set symmetrically along the diagonal. After the Transformer-Encoder processes the input representations (i.e., $\{r_{i}|i\in[1,8]\}_{v}$, $\{P_{9m}|m\in[1,M]\}_{v}$ and $L$ learnable vectors) and generates the output, we extract and save those $L$ optimizable vectors from the output. These $L$ (processed) output learnable vectors are then individually processed by an MLP into vectors of double length. These $L$ doubled-length vectors are transformed into $L$ vector using the reparameterization technique[8]. The process of reparameterization, which is a practical approach used in VAE[8], can be visualized as Figure 14.
@@
These $L$ reparameterized vector are regarded as the representations of $L$ auxiliary samples. It is worth noting that $L$ is a hyperparameter that can be freely set, determining how many judicious auxiliary samples Tine generates for Valen in one instance. In this paper, $L$ is set to 9.
In summary, our Tine method is proposed after summarizing the working principles of probability-highlighting solvers such as Valen. We believe that such solvers establish a solution distribution for each instance of a reasoning problem, with the optimization objective being to assign high probability to primary samples and low probability to auxiliary samples. This means that the optimization target of the solution distribution is not aligned with the distribution of correct solutions to the reasoning problems, but rather with the distribution delimited by both primary and auxiliary samples. Unfortunately, this delimited distribution is often much coarser than the correct solution distribution. To make the distribution delimited by both primary and auxiliary samples closer to the correct solution distribution, we designed Tine based on adversarial learning. Although subsequent experiments demonstrated that Tine can effectively improve Valen’s performance, we believe that the Tine method still fails to address the underlying essence of the problem. Therefore, further exploration is conducted in this paper.
@@
Based on the analysis in the previous section, auxiliary samples provided in reasoning tasks constrain the standard deviation of the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$, while primary samples locate the mean of the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$. The Tine method generates more judicious auxiliary samples to further constrain the standard deviation. Although experimental results demonstrate the effectiveness of Tine, it remains uncertain whether Tine constrains the standard deviation too little or not enough. Furthermore, Tine based on adversarial learning also encounters all the issues associated with adversarial learning, such as unstable training processes and difficulties in determining the completeness of training. Given this series of challenges, this paper proposes the Funny (Framework Utilizing Neural Networks for Yielding Image Representation Distribution) method.
@@
Our Funny method argues that instead of laboriously adopting adversarial learning strategies to make the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ provided by the solver closer to the correct solution distribution, it is more effective to directly specify the range of correct solutions or, alternatively, to explicitly define the details of the correct solution distribution. This paper will next analyze the definition of correct solutions for image progressive pattern reasoning problems such as RPM and image clustering reasoning problems such as Bongard-Logo.

### v3

@@ IV-B Valen targeting Bongard-Logo problem
We regard Valen as a strong baseline for image progression pattern reasoning and as a model capable of tackling image clustering reasoning problem such as Bongard-Logo. To adapt Valen to Bongard-Logo, we transform the task accordingly.
In a given Bongard-Logo instance, we denote the 14 images as $\{x_{i}|i\in[1,14]\}$. Specifically, the images in the primary group are $\{x_{i}|i\in[1,6]\}$, while the images in the auxiliary group are $\{x_{i}|i\in[8,13]\}$. Furthermore, $x_{7}$ is to be assigned to the primary group, and $x_{14}$ is to be assigned to the auxiliary group.
This paper proposes that image clustering reasoning problems, such as Bongard-Logo, can be approached from different perspectives. Specifically, these tasks can be recast as option-evaluation problems, analogous to those in the RPM task. To elaborate, the images in the primary group $\{x_{i}|i\in[1,6]\}$ can be regarded as a problem statement akin to that in RPM, while the test images and those in the auxiliary group $\{x_{i}|i\in[7,14]\}$ act as an option pool analogous to the one in RPM. This transformation is shown in Figure 8.
@@
After this transformation, Valen can be effectively applied to image clustering reasoning problems such as Bongard-Logo.
It is noteworthy that when transferring Valen to Bongard-Logo, some adjustments are still necessary to accommodate the data characteristics of Bongard-Logo. Specifically, due to the scarcity of training instances in the Bongard-Logo problem, the ViT-based image representation extractor may struggle to perform effectively [21, 32]. Furthermore, although we transform the Bongard-Logo problem into an RPM-like format, it is unnecessary to analyze the progressive patterns among Bongard-Logo images, as such patterns do not exist in clustering reasoning problems like Bongard-Logo.
Therefore, we have redesigned the image representation extractor, originally illustrated in Figure 4, into a convolutional network concatenated with a Transformer-Encoder. The new design is illustrated in Figure 9.
@@
In this figure, when a Bongard-Logo image is processed by Valen’s new feature extractor, it is first encoded into a feature map by ResNet18. This map is then decomposed into multiple tokens and fed into a Transformer-Encoder to obtain multi-viewpoint representations. The design combines CNN’s sample efficiency with ViT’s multi-viewpoint modeling capacity [1, 2, 21, 32, 27] and has been used in previous work [27].
@@
In this figure, we abandon the use of MLP to extract the row and column progressive information from the incomplete matrix in Bongard-Logo, because the transformed Bongard-Logo problem only contains clustering patterns and not progressive patterns. We then feed the representations of all images within the Bongard-Logo incomplete matrix as tokens into the Transformer-Encoder to extract $M$ clustering patterns. When processing Bongard-Logo, $M$ is set to 6.
@@ V Tine, a Method for Advancing the Solution Distribution Boundary
Current end-to-end visual abstract reasoning solvers, including Valen, are keen on assigning probabilities to each option. Although their loss functions are diverse and their backbone architectures vary significantly, most adhere to this paradigm. We term such solvers “probability-highlighting solvers”.
@@
Therefore, we can interpret these solvers as a conditional probability distribution governing the solutions to abstract reasoning problems. Taking RPM solvers as examples, they can be abstracted as a conditional distribution, denoted as $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$. (Accordingly, when Valen is applied to Bongard-Logo problems, it can be abstracted as $P(x_{\alpha}|\{x_{i}\}_{i=1}^{6};\theta)$). Here, $\{x_{i}\}_{i=1}^{8}$ is the statement of the reasoning problem, $\theta$ represents the learnable parameters of the solver, and $x_{\alpha}$ refers to the option that is pending verification. This distribution gives the probability that the solver’s parameters $\theta$ deem $x_{\alpha}$ to be the solution to the context $\{x_{i}\}_{i=1}^{8}$.
@@
The optimization of a probability-highlighting solver corresponds to tuning the parameters $\theta$ within the conditional distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$. Although a wide range of loss functions can be employed, they all share the same objective: adjust parameters $\theta$ so that the distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ assigns high probability to the primary samples and low probability to the auxiliary samples. Such an optimization objective makes us realize that the optimization process of the probability-highlighting solver is not aimed at fitting the distribution of correct solutions to visual abstract reasoning problems, but rather is dedicated to fitting a distribution jointly delimited by primary and auxiliary samples.
@@
To simplify the explanation of our findings, we assume that the conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ follows a Gaussian distribution, so we write it explicitly as $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\mu_{\theta},\sigma_{\theta})$, where $\mu_{\theta}$, $\sigma_{\theta}$ represent the mean and standard deviation, respectively. Then, the negative logarithm of this Gaussian density can be expressed as follows:
$-\log(P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\mu_{\theta},\sigma_{\theta}))=\frac{{(x_{\alpha}-\mu_{\theta})^{2}}}{{2\sigma^{2}_{\theta}}}-\log(2\pi|\,\sigma_{\theta}|\,)$ (1)
@@
Accordingly, the optimization objective is to tune the values of $\mu_{\theta}$ and $\sigma_{\theta}$ so that when primary samples (i.e., the correct RPM options or samples within the primary group in Bongard-Logo) are plugged into formula (1), the formula returns a smaller value; conversely, when auxiliary samples (i.e., the incorrect RPM options or the samples within the auxiliary group in a Bongard-Logo problem) are plugged into formula (1), the formula returns a larger value.
@@
Evidently, minimizing the return value of the formula (1) amounts to minimizing the Euclidean distance between $\mu_{\theta}$ and $x_{\alpha}$; maximizing it corresponds to minimizing the value of $\sigma_{\theta}$. In other words, the probability-highlighting solver positions the mean of the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ for the reasoning problem through primary samples, while the standard deviation of the solution distribution is constrained by auxiliary samples.
@@
It appears to be a discouraging reality that the learning objective of a probability-highlighting solver is not to determine the distribution of the correct solutions to reasoning problems, but rather to approximate a distribution that is delimited by both primary and auxiliary samples within those problems. This suggests that the reasoning accuracy of the probability-highlighting solver is constrained. The standard deviation of the distribution delimited by the auxiliary samples is necessarily greater than or equal to the standard deviation of the correct solution distribution. In other words, the distribution delimited by both primary and auxiliary samples is often much coarser than the distribution of correct solutions. Aiming to approximate the former as the optimization objective is akin to “aiming for mediocrity and achieving even less”. This discouraging reality can be visualized in Figure 11. Note that this figure is intended to schematically represent our analysis results, rather than to depict the actual form of the solution distribution or the correct solution distribution.
@@
This paper argues that to further improve the reasoning accuracy of Valen in solving RPM problems and Bongard-Logo problems, it is necessary to provide Valen with more judicious auxiliary samples. By “more judicious auxiliary samples”, we refer to those that can better delimit the solution distribution provided by Valen. These more judicious auxiliary samples have the potential to narrow down the standard deviation of the solution distribution. The visualization of these more judicious auxiliary samples can be found in Figure 12.
@@
As shown in this figure, the judicious auxiliary samples we desire appear more authentic or closer to primary samples from Valen’s perspective than the original auxiliary samples. However, since judiciousness is evaluated from Valen’s perspective, it lacks human interpretability; samples with smaller or fewer attribute shifts are not necessarily deemed more judicious by Valen.
To this end, this paper argues that we can use adversarial learning to supply Valen with auxiliary samples it deems more judicious. Given this insight, we design the Tine (Technique for edge Identification of solution distribution based on adversarial Networks) method. Specifically, we adopt adversarial generation techniques similar to GAN [9] or WGAN [39] to generate more judicious samples for Valen. However, unlike GAN or WGAN, whose objective is to train robust and high-performance generators, we use adversarial learning to refine Valen, which acts as the discriminator.
Our objective is to design an auxiliary sample generator for Valen. Its task is to produce samples that receive higher probabilities from Valen than the original auxiliary samples. Accordingly, Valen’s role is to classify these generated samples as low-scoring ones. During training, the parameters of Valen and the generator are updated alternately and exclusively, undergoing standard adversarial refinement.
Given that our goal is to generate more judicious auxiliary samples from Valen’s perspective, we are not concerned with their interpretability, attribute values, or whether they strictly conform to the nature of reasoning problems. Consequently, rather than forcing the generator to produce pixel-level auxiliary samples, which would incur greater computational overhead and optimization difficulties, we opt to generate representations of judicious auxiliary samples. By outputting only multi-viewpoint representations, the generator enjoys a simpler architecture and becomes easier to optimize. The generator is shown in Figure 13. Since we have already transformed the Bongard-Logo problem into an RPM-like problem, we will not further elaborate on the structure of the generator for the Bongard-Logo version.
@@
As depicted in Figure 13, the left gray dashed box encloses the Valen model, while the right gray dashed box encloses the generator designed for Valen. The backbone of this generator is a standard Transformer-Encoder, which generates representations of more judicious auxiliary samples based on the representations of the abstract reasoning problem statement $\{r_{i}|i\in[1,8]\}_{v}$, their progressive patterns $\{P_{9m}|m\in[1,M]\}_{v}$, and $L$ learnable vectors. Both the representations and the progressive patterns of the statement are provided by Valen, while the $L$ learnable vectors must be additionally introduced for the generator. It is noteworthy that when dealing with PGM problems within RPM tasks, we have designed special positional encoding for the representations of the problem statement. This positional encoding can be specifically attached to incomplete matrices with missing option positions and is set symmetrically along the diagonal. After the Transformer-Encoder processes the input representations (i.e., $\{r_{i}|i\in[1,8]\}_{v}$, $\{P_{9m}|m\in[1,M]\}_{v}$ and $L$ learnable vectors) and generates the output, we extract and save those $L$ learnable vectors. These processed vectors then pass through an MLP that doubles their dimension. The resulting $L$ doubled-length vectors are transformed back into $L$ vectors using the reparameterization technique [10]. The details of this technique are shown in Figure 14.
@@
These $L$ reparameterized vectors are regarded as the representations of $L$ auxiliary samples. It is worth noting that $L$ is a hyperparameter that can be freely set, determining how many judicious auxiliary samples Tine generates for Valen in one instance. Here, $L$ is set to 9.
In short, our Tine method was conceived after summarizing the working principles of probability-highlighting solvers. We contend that such solvers establish a solution distribution for every instance of a reasoning problem and adjust it so that primary samples receive high probability while auxiliary samples receive low probability. This means their optimization target is not to align with the distribution of correct solutions to the reasoning problems, but rather to align with the distribution jointly delimited by primary and auxiliary samples. Unfortunately, this delimited distribution is often much coarser than the correct one. To bring the former closer to the latter, we designed Tine based on adversarial learning.
@@
Based on the analysis in the previous section, auxiliary samples provided in reasoning tasks constrain the standard deviation of the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$, while primary samples locate the mean of the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$. The Tine method generates more judicious auxiliary samples to further constrain the standard deviation. However, it remains unclear whether this constraint is insufficient or excessive. Moreover, being rooted in adversarial learning, Tine inevitably inherits the associated drawbacks, such as unstable training and difficulty in assessing convergence. Faced with these challenges, we propose the Funny (Framework Utilizing Neural Networks for Yielding Image Representation Distribution) method.
@@
Our Funny method posits that, rather than laboriously adopting adversarial learning to make the solver’s distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ close to the correct solution distribution, it is more effective to directly specify the range of correct solutions or to explicitly characterize the correct solution distribution. Toward this goal, the present section dissects the correct solutions for both image progression pattern and image clustering reasoning problems.
VI-A Key attributes and non-key attributes